# supplement.pdf

### v1

@@ Abstract
Note that the observable ${\mathcal{O}}_{\rm rev}$ here is independent of $p$, in contrast to ${\mathcal{O}}$ in Theorem 1 that is a function of $p$. We sketch the idea of Theorem 2 and leave the full proof to SM . Note that all approximations $\approx$” below stem from the light-cone approximation in Eq. (8).
Focusing on $d=2$ (generalization is straightforward) as shown in Fig. 2 and recalling that $|\tilde{\theta}|\sim L^{-2}$ and $T\sim L$, we first expand in the noiseless case,
$\langle\widetilde{P}_{I}\rangle_{\rho_{\theta}^{(0)}}$ $=\bra{\bm{0}}\mathrm{e}^{\mathrm{i}\tilde{\theta}Z(T)}P_{\widetilde{I}}\mathrm{e}^{-\mathrm{i}\tilde{\theta}Z(T)}\ket{\bm{0}}$
$=1-\frac{\tilde{\theta}^{2}}{2}\bra{\bm{0}}[Z(T),[Z(T),P_{\widetilde{I}}]]\ket{\bm{0}}+\cdots.$ (13)
The first-order term vanishes similarly as Eq. (11), and we show that the second-order term is $\sim\tilde{\theta}^{2}L^{4}$. (Note that we have required that $|\tilde{\theta}|$ is sufficiently large to guarantee that the second-order derivative is dominant.) We expand $Z=\sum_{J}Z_{J}$ and observe that only $J=I$ or neighbors of $I$ contribute, because otherwise $[Z_{J}(T),P_{\widetilde{I}}]=0$ since they do not overlap in space from Eq. (8). Furthermore, when $L/T$ is a sufficiently large constant, the coefficient is dominated by the term
$\bra{\bm{0}}[Z_{I}(T),[Z_{I}(T),P_{\widetilde{I}}]]\ket{\bm{0}}\\ \approx 2\,C_{\rho}(Z_{I},Z_{I})\geq 2c_{Z}L^{2}T^{2},$ (14)
which can be verified from direct computation and Eq. (10). The reason is that the sum of all other terms is $\lesssim LT^{3}$, which becomes subdominant when $L/T$ is sufficiently large. To show this, we note that for $i\in I,j\in J\neq I$,
$\bra{\bm{0}}[Z_{i}(T),[Z_{j}(T),P_{\widetilde{I}}]]\ket{\bm{0}}\approx 0,$ (15)
unless the two sites are close ($\mathrm{d}(i,j)\leq 2T$), as shown in Fig. 2. Then the sum of Eq. (15) over $i,j$ is bounded by $\sum_{j\text{ close to }I}\sum_{i\text{ close to }j}\lesssim LT\cdot T^{2}$.
Above we showed, at sufficiently large $L/T$, the second-order coefficient in Eq. (Small correlation is sufficient for optimal noisy quantum metrology) is $\sim L^{4}$. This guarantees $-\partial_{\theta}\langle\widetilde{P}_{I}\rangle_{\rho_{\theta}^{(0)}}\sim\tilde{\theta}\times\text{[the second-order coefficient]}\sim L^{2}$, and thus
$\partial_{\theta}\left\langle{\mathcal{O}}_{\rm rev}\right\rangle_{\rho_{\theta}^{(0)}}\sim N.$ (16)
(Note that the higher orders of $\tilde{\theta}$ in Eq. (Small correlation is sufficient for optimal noisy quantum metrology) are safely ignored by similar locality analysis.) Since faraway groups are hardly correlated,
$\left\langle\Delta{\mathcal{O}}^{2}_{\rm rev}\right\rangle_{\rho_{\theta}^{(0)}}\sim\sum_{I}\big{\langle}\Delta\widetilde{P}_{I}^{2}\big{\rangle}_{\rho_{\theta}^{(0)}}\lesssim N/L^{2},$ (17)
leading to Eq. (12) for the noiseless case.
Finally, we consider the noisy case $p>0$. Eq. (17) still holds (where $\rho_{\theta}^{(0)}$ is replaced with $\rho_{\theta}$) because adding noise does not create long-range correlation Poulin (2010). It then remains to show Eq. (16) is also robust. For illustrative purposes, we focus on dephasing noise here, i.e., $\mathcal{N}_{\theta,p}={\mathcal{D}}_{\theta,p}$, yielding
$\left\lvert\partial_{\theta}\big{(}\big{\langle}{\widetilde{P}_{I}}\big{\rangle}_{\rho_{\theta}}-\big{\langle}{\widetilde{P}_{I}}\big{\rangle}_{\rho_{\theta}^{(0)}}\big{)}\right\rvert=\left\lvert\big{\langle}{\big{[}Z,{\mathcal{D}}_{0,p}^{\otimes N}(\widetilde{P}_{I})-\widetilde{P}_{I}}\big{]}\big{\rangle}_{\rho_{\theta}^{(0)}}\right\rvert$
$\lesssim L^{2}\left\lVert{\mathcal{D}}_{0,p}^{\otimes c_{0}L^{2}}(\widetilde{P}_{I})-\widetilde{P}_{I}\right\rVert\lesssim L^{4}p,$ (18)
where we recall that the dual map of the dephasing noise ${\mathcal{D}}_{0,p}$ is itself, and use the locality of $\widetilde{P}_{I}$ to restrict the channel to a region of size $c_{0}L^{2}$ with $c_{0}\sim 1$. Taking $pL^{2}\leq c_{\rm no}$ and $c_{\rm no}$ sufficiently small guarantees the contribution from noise to Eq. (16) is subdominant relative to the noiseless result.
Quantum domino dynamics with local measurement.— We have shown that any state generated by local dynamics with correlations of the form Eq. (10) is metrologically optimal when $L^{d}\sim 1/p$. Here we present a simple application of this framework using quantum domino dynamics Lee and Khitrin (2005); Yoshinaga et al. (2021).
@@
which can be engineered as the effective Hamiltonian for the transerve-field Ising model at weak $X$ field Lee and Khitrin (2005). The vacuum |$\ket{\bm{0}}$ is an eigenstate of $H_{\rm do}$, and, more generally, the number of domain walls (DWs)—transitions from |$\ket{0}$ to |$\ket{1}$ or vice versa—is conserved. If the system starts from the initial state |$\ket{1_{1}}:=\ket{10\cdots 0}$, it only traverses the state space with one DW $\mathrm{e}^{-\mathrm{i}TH_{\rm do}}\ket{1_{1}}=\sum_{i=1}^{N}a_{i}\ket{1_{i}}$, where |$\ket{1_{i}}$ is the state with $1$ on the first $i$ qubits and 00 on the rest, and where the coefficients $a_{i}$ are determined by $T$. This single-particle hopping process of the DW justifies the name “domino”. Before $T\gtrsim N$ (i.e., when the DW propagates to the right), $a_{i}$ is centered around $i=vT$ with a velocity $v$. Then, an initial state |$\ket{+}_{1}\otimes|\bm{0}\rangle$ would evolve to
@@
Figure 3: Optimal noisy metrology for $N=200$ qubits under dephasing, using partitioned domino state Eq. (21) generated from quantum domino dynamics, and local measurements ${\mathcal{O}}_{\rm do}$ of the form Eq. (25). The $y$-axis quantifies $(\delta\theta)^{-2}/N$ for local estimation at $\theta\approx 0$. For each solid line, the spin chain is divided to $N/L$ groups each of size $L$, and the generating time $T\sim L$ is suitably chosen (see SM for details). For any noise probability $p\ll 1$ in Eq. (1), by choosing suitable $T$ and $L$, the performance approaches the bound Eq. (2) (pink dashed line) up to a factor $\approx 2$.
@@
|$\ket{\Psi_{\rm do}}:=\mathrm{e}^{-\mathrm{i}TH_{\rm do}}\left(\ket{+}_{L}\otimes\ket{+}_{2L}\otimes\cdots\right)\otimes|\bm{0}\rangle_{\rm rest}.$ (21)
This final state, which we call the partitioned domino state, satisfies Eq. (10) for $T\lesssim L$, because the two DWs of each $\ket{+}$ roughly propagate in opposite directions with constant velocity. We numerically study noisy metrology using |$\ket{\Psi_{\rm do}}$ on a chain of $200$ qubits, and Fig. 3 shows the results. Remarkably, this protocol is not only optimal asymptotically, but also quantitatively. By adjusting $L$ for different $p$, the precision $(\delta\theta)^{2}$ is only roughly twice larger than the ultimate limit Eq. (2).
Besides having nearly the same metrological power as the quantum parity state, the domino-state protocol requires only local measurements without any time-reversed evolution. To show this, we focus first on the simplest one-DW case with state Eq. (20). In SM , we prove the following for various combinations $\mathcal{N}_{\theta,p}$ of $\theta$-imparting unitary evolution and a general noisy channel.
@@
In a rotated local basis (i.e., redefining states |$\ket{1}_{i}$ with extra phases) such that $\mathrm{Arg}(a_{i})=-\frac{\pi}{2}(1+i)-2\theta_{\rm pr}i$, define the measurement observable
${\mathcal{O}}_{\rm do}:=\sum_{vT/2\leq i\leq 2vT}|a_{i}|\mathbf{Y}_{i},$ (22)
where $\mathbf{Y}_{i}:=\bigotimes_{j=1}^{i}Y_{j}$. If
@@
where $\rho_{\theta}=\mathcal{N}_{\theta,p}^{\otimes N}(\ket{\psi_{\rm do}}\bra{\psi_{\rm do}})$ with noise $p\leq c_{\rm do}/(vT)$ for a constant $c_{\rm do}$.
Eq. (22) can be realized by $Y$-basis local measurements with efficient classical post-processing. The condition Eq. (23) means that $|a_{i}|^{2}$, as a probability distribution, is concentrated around $vT$, which can be verified from the exact solution (Bessel function) of $a_{i}$ Roubert et al. (2010). The idea is as follows. |$\ket{\psi_{\rm do}}$ is roughly a superposition of GHZ states |$\ket{\mathrm{GHZ}}_{1\cdots i}\otimes\ket{\bm{0}}$ of different size $i$. For each GHZ state, the canonical observable is $\mathbf{Y}_{i}$. The problem is that each such observable does not have large sensitivity, i.e., $\partial_{\theta}\left\langle\mathbf{Y}_{i}\right\rangle\sim|a_{i}|\ll 1$, because the DW wave packet spreads. To overcome this, we use a suitable linear combination of them, namely, ${\mathcal{O}}_{\rm do}$, to accumulate the small derivatives into $\partial_{\theta}{\mathcal{O}}_{\rm do}\sim\sum_{i}|a_{i}|^{2}\sim 1$. It turns out this linear combination does not lead to large variance of ${\mathcal{O}}_{\rm do}$ due to the lack of interference among the GHZ states in |$\ket{\psi_{\rm do}}$, making ${\mathcal{O}}_{\rm do}$ a suitable observable.
@@
$\bigg{(}\sum_{i,j}b_{i,j}\mathbf{Y}_{i,j}\bigg{)}\otimes\bigg{(}\sum_{i,j}b_{i,j}\mathbf{Y}_{L+i,L+j}\bigg{)}\otimes\cdots,$ (25)
where the real coefficients $b_{i,j}$ come from simulating the two-DW subspace, which only requires polynomial (in $L$) resource of classical computation. See SM for details.
Spin-squeezing.— We show in SM that certain types of spin-squeezed states (SSS) Kitagawa and Ueda (1993); Ma et al. (2011), though demanding to prepare Block et al. (2023), are another example of metrological optimality. Although spin squeezing is beyond the regime of Theorem 1, certain boundedness of correlation functions is also sufficient, suggesting the general usefulness of correlation bounds for noisy metrology.
@@
Beyond state optimality, we also address the problem of efficient measurement schemes for metrology. In particular, the domino example suggests that local measurements apply to wider sensing scenarios beyond GHZ: Although a local measurement scheme that exactly saturates the QFI Liu et al. (2023) may not exist, the QFI can still be saturated up to a moderate prefactor. The example also raises an interesting connection between quantum metrology and many-body dynamics, hinting at the possibility of other natural (e.g. time-independent) Hamiltonian dynamics generating high-QFI states in higher spatial dimensions.

### v2

@@ Abstract
Theorem 2 implies locally generated optimal states, combined with the time-reversal measurement protocol, achieve the optimal QFI when choosing $M\sim 1/p$. An additional advantage of the time-reversal measurement protocol is the observable ${\mathcal{O}}_{\rm rev}$ is independent of $p$, in contrast to ${\mathcal{O}}$ in Theorem 1 that may be a function of $p$. A proof sketch can be found in End Matter, with a full proof in SM .
Quantum domino dynamics with on-site measurement.— We have shown that any state generated by local dynamics with correlations of the form Eq. (10) is metrologically optimal when $L^{d}\sim 1/p$. Here we present an explicit example of this framework using quantum domino dynamics Lee and Khitrin (2005); Yoshinaga et al. (2021), where an on-site measurement protocol without time-reversal is sufficient to achieve the optimal QFI.
@@
which can be engineered as the effective Hamiltonian for the transerve-field Ising model at weak $X$ field Lee and Khitrin (2005). The vacuum |$\ket{\bm{0}}$ is an eigenstate of $H_{\rm do}$, and, more generally, the number of domain walls (DWs)—transitions from |$\ket{0}$ to |$\ket{1}$ or vice versa—is conserved. If the system starts from the initial state |$\ket{1_{1}}:=\ket{10\cdots 0}$, it only traverses the state space with one DW $\mathrm{e}^{-\mathrm{i}TH_{\rm do}}\ket{1_{1}}=\sum_{i=1}^{N}a_{i}\ket{1_{i}}$, where |$\ket{1_{i}}$ is the state with $1$ on the first $i$ qubits and 00 on the rest and the coefficients $a_{i}$ are determined by $T$. This single-particle hopping process of the DW justifies the name “domino”. Before $vT\gtrsim N$ (i.e., when the DW propagates to the right), $a_{i}$ is centered around $i=vT$ with a velocity $v$. Then, an initial state |$\ket{+}_{1}\otimes|\bm{0}\rangle$ would evolve to
@@
Figure 2: Optimal noisy metrology for $N=200$ qubits under dephasing, using partitioned domino state Eq. (16) generated from quantum domino dynamics, and on-site measurements ${\mathcal{O}}_{\rm do}$ of the form Eq. (20). The $y$-axis quantifies $(\delta\theta)^{-2}/N$ for local estimation at $\theta=0$. For each solid line, the spin chain is divided to $N/L$ groups each of size $L$, and the generating time $T\sim L$ is suitably chosen (see SM for details). For any noise probability $p\ll 1$ in Eq. (1), by choosing suitable $T$ and $L$, the performance approaches the bound in Eq. (2) (pink dashed line) up to a factor $\approx 2$.
@@
|$\ket{\Psi_{\rm do}}:=\mathrm{e}^{-\mathrm{i}TH_{\rm do}}\left(\ket{+}_{1}\otimes\ket{+}_{L+1}\otimes\cdots\right)\otimes|\bm{0}\rangle_{\rm rest}.$ (16)
This final state, which we call the partitioned domino state, satisfies Eq. (10) for $T\lesssim L$, because the two DWs of each $\ket{+}$ roughly propagate in opposite directions with constant velocity. We numerically study noisy metrology using |$\ket{\Psi_{\rm do}}$ on a chain of $200$ qubits, and Fig. 2 shows the results. Remarkably, this protocol is not only optimal asymptotically, but also quantitatively. By adjusting $L$ for different $p$, the precision $(\delta\theta)^{2}$ is only roughly twice larger than the known lower bound Eq. (2).
Besides having the optimal metrological power, the domino-state protocol requires only on-site measurements without any time-reversed evolution. To show this, we focus first on the simplest one-DW case with state Eq. (15) and prove the following in SM :
@@
In a rotated on-site basis (i.e., redefining states |$\ket{1}_{i}$ with extra phases) such that $\mathrm{Arg}(a_{i})=-\frac{\pi}{2}(1+i)-2\theta_{\rm pr}i$, define the measurement observable
${\mathcal{O}}_{\rm do}:=\sum_{vT/2\leq i\leq 2vT}|a_{i}|\mathbf{Y}_{1,i},$ (17)
where $\mathbf{Y}_{1,i}:=\bigotimes_{j=1}^{i}Y_{j}$. If
@@
if noise $p\leq c_{\rm do}/(vT)$ for a constant $c_{\rm do}$.
Eq. (17) can be realized by $Y$-basis on-site measurements with efficient classical post-processing. The condition Eq. (18) means that $|a_{i}|^{2}$, as a probability distribution, is concentrated around $vT$, which holds before $vT\gtrsim N$.
The idea to choose Eq. (17) as the observable is as follows. |$\ket{\psi_{\rm do}}$ is roughly a superposition of GHZ states |$\ket{\mathrm{GHZ}}_{1\cdots i}\otimes\ket{\bm{0}}$ of different size $i$. For each GHZ state, $\mathbf{Y}_{1,i}$ is an optimal observable to sense the $Z$-axis magnetic field. However, each such observable does not have large sensitivity, i.e., $\partial_{\theta}\left\langle\mathbf{Y}_{1,i}\right\rangle\sim|a_{i}|i\ll i$, because the DW wave packet spreads. To overcome this, we use a suitable linear combination of them, namely, ${\mathcal{O}}_{\rm do}$, to accumulate the small derivatives into $\partial_{\theta}{\mathcal{O}}_{\rm do}\sim\sum_{i}|a_{i}|^{2}\sim 1$. It turns out this linear combination does not lead to large variance of ${\mathcal{O}}_{\rm do}$ due to the lack of interference among the GHZ states in |$\ket{\psi_{\rm do}}$, making ${\mathcal{O}}_{\rm do}$ a suitable observable.
@@
$\sum_{\text{integer }I}\,\sum_{1-L/2\leq i\leq j\leq L/2}b_{i,j}\mathbf{Y}_{IL+i,IL+j}~{},$ (20)
where the real coefficients $b_{i,j}$ come from simulating the two-DW subspace for the group $I=0$, which only requires polynomial (in $L$) resource of classical computation. See SM for details.
Spin-squeezing.— We have established a sufficient condition for metrological optimality. Here we show that this small-correlation condition (Eq. (6)) is not necessary: spin-squeezed states (SSS) Kitagawa and Ueda (1993); Ma et al. (2011) do not satisfy it in general, but are optimal by the following theorem proven in SM , which generalizes special SSSs and noise analyzed in Ulam-Orgikh and Kitagawa (2001); Zhou and Jiang (2021); Brask et al. (2015).
Theorem 4.
Suppose state $\rho$ satisfies
$\left\langle X\right\rangle_{\rho}$ $=\Theta(N),\quad\left\langle\Delta Y^{2}\right\rangle_{\rho}=N\xi^{2},$ (21)
$\left\langle\Delta S^{2}\right\rangle_{\rho}$ $=\mathrm{O}(N/\xi^{2}),\quad\forall S:=\textstyle\sum_{i}S_{i}\text{ with }\left\lVert S_{i}\right\rVert=1$ (22)
where $X=\sum_{i}X_{i}$ etc and $S_{i}$ denotes an operator on $i$. Its metrological power is robust in noisy metrology with $p=\mathrm{O}(\xi^{2})$:
$\mathcal{F}(\rho_{\theta})\geq\frac{1}{\left\langle\Delta Y^{2}\right\rangle_{\rho_{\theta}}}\left(\partial_{\theta}\left\langle Y\right\rangle_{\rho_{\theta}}\right)^{2}\Big{|}_{\theta=0}=\Theta(N/\xi^{2}).$ (23)
Eq. (21) is a standard definition of SSS: $\rho$ is polarized in the $X$ direction and squeezed by a factor $\xi\ll 1$ in the $Y$ direction. For example, extremely squeezed states with $\xi=\Theta(1/\sqrt{N})$ exist in the permutation-symmetric subspace of the $N$ qubits, for which any two groups of qubits have a constant correlation $C_{\rho}(I;J)=\Theta(1)$, violating Eq. (6). Explanations on the violation of Eq. (6) for any sufficiently small $\xi$ can be found in SM . Although beyond the scope of Theorem 1, such SSSs are optimal by Theorem 4, as long as its “polarization” $S$ along any direction in Eq. (22) is at most stretched by a factor $\mathrm{O}(1/\xi)$, the minimal amount given by the uncertainty principle (as one can see by choosing $S=Z$). The intuition here is that under general noise, the metrological observable $Y$ may mix with “orthogonal” directions $S$ and thus increase its variance; Eq. (22) ensures this mixing is weak enough. The condition Eq. (22) suggests the general usefulness of correlation bounds for noisy metrology.
Eq. (23) shows that on-site measurements suffice for noisy metrology with SSS. However, it is not straightforward to prepare SSSs from natural (i.e., time-independent) local dynamics: The squeezing parameter $\xi$ has a limited scaling with $N$ in recent proposals Block et al. (2024); Comparin et al. (2022); Roscilde et al. (2024), and it is argued that the 1d case can only generate $\xi\sim 1$ independent of $N$. This is in sharp contrast with the preparable domino state above, which has optimal robustness against noise and on-site measurement schemes. Nevertheless, it is unclear whether the domino dynamics can be generalized to higher dimensions. The small-correlation states and SSSs are thus complementary classes of metrological resources, with each offering advantages depending on the context.
@@
Beyond state optimality, we also address the problem of efficient measurement schemes for metrology. In particular, the domino example suggests that on-site measurements apply to wider sensing scenarios beyond GHZ: Although an on-site measurement scheme that exactly saturates the QFI Liu et al. (2023) may not exist, the QFI can still be saturated up to a moderate prefactor.